OpenStudy (anonymous):

Frank is four years older than bill. Six years ago, he was twice as old as Bill. how old is frank now?

OpenStudy (vishweshshrimali5):

Let, the age of Frank is f and that of Bill is b (at present).

OpenStudy (vishweshshrimali5):

Now, it is given that Frank is 4 years older than Bill, so, f = 4b

OpenStudy (vishweshshrimali5):

Ohh sorry that was f = 4 + b

OpenStudy (anonymous):

ok

OpenStudy (vishweshshrimali5):

Good , now 6 years ago what would have been the ages of Frank and Bill ?

OpenStudy (vishweshshrimali5):

They would be: Age of Frank = f - 6 Age of Bill = b - 6 Got it ?

OpenStudy (anonymous):

ya i am just bad at like substituting and eliminating

OpenStudy (vishweshshrimali5):

Its okay.

OpenStudy (vishweshshrimali5):

Now the question says: "Six years ago, he(that is Frank) was twice as old as Bill"

OpenStudy (anonymous):

f-6=2(b-6)

OpenStudy (vishweshshrimali5):

Very nice.

OpenStudy (vishweshshrimali5):

Now can you solve this?

OpenStudy (anonymous):

thats the part I'm stuck on

OpenStudy (vishweshshrimali5):

Its ok

OpenStudy (vishweshshrimali5):

(f-6) = 2(b-6) => f - 6 = 2b - 2*6 => f - 6 = 2b - 12 => f = 2b - 12 + 6 => f = 2b - 6 Do you get this part ?

OpenStudy (anonymous):

ooooo okkk i see wat i got wrong thank u!!

OpenStudy (vishweshshrimali5):

No problem.

OpenStudy (vishweshshrimali5):

Now, we have two equations : f = 4 + b and f = 2b - 6 Now, lets see if we can solve them.

OpenStudy (vishweshshrimali5):

What I am going to do is , I am going to take value of "f" from 1st equation and going to put it in 2nd equation. Okay ?

OpenStudy (vishweshshrimali5):

@dbaseball You there ?